OpenStudy (imsleeziboii):

simplify (-3x^-4)^2 x

OpenStudy (freckles):

use this law of exponents \[(r s^y)^a=(r^1 s^y)^a=r^{1 \cdot a} s^{y \cdot a}\]

OpenStudy (imsleeziboii):

ok

OpenStudy (freckles):

or you can do this \[(-3x^{-4})^2=(-3x^{-4})(-3x^{-4})=(-3)(-3)x^{-4}x^{-4}\]

OpenStudy (imsleeziboii):

im on a different problem now

OpenStudy (freckles):

k